# Pushing the frontiers of subexponential FPT time for Feedback Vertex Set

### v1

@@ Abstract.
The paper deals with the Feedback Vertex Set problem parameterized by the solution size. Given a graph $G$ and a parameter $k$, one has to decide if there is a set $S$ of at most $k$ vertices such that $G-S$ is acyclic. Assuming the Exponential Time Hypothesis, it is known that FVS cannot be solved in time $2^{o(k)}n^{\mathcal{O}(1)}$ in general graphs. To overcome this, many recent results considered FVS restricted to particular intersection graph classes and provided such $2^{o(k)}n^{\mathcal{O}(1)}$ algorithms.
In this paper we provide generic conditions on a graph class for the existence of an algorithm solving FVS in subexponential FPT time, i.e. time $2^{k^{\varepsilon}}\mathop{\rm poly}(n)$, for some $\varepsilon<1$, where $n$ denotes the number of vertices of the instance and $k$ the parameter. On the one hand this result unifies algorithms that have been proposed over the years for several graph classes such as planar graphs, map graphs, unit-disk graphs, pseudo-disk graphs, and string graphs of bounded edge-degree. On the other hand it extends the tractability horizon of FVS to new classes that are not amenable to previously used techniques, in particular intersection graphs of “thin” objects like segment graphs or more generally $s$-string graphs.
@@ 1.1. Context
Given an $n$-vertex graph $G$ and a parameter $k\in\mathbb{N}$, the Feedback Vertex Set problem (FVS for short) asks whether there exists a set $S$ of at most $k$ vertices such that $G-S$ has no cycle. This is a fundamental decision problem in graph theory and one of Karp’s 21 NP-complete problems. Because of its hardness in a classical setting, the problem has been widely studied within the realm of parameterized complexity. This line of research aims to investigate the existence of FPT algorithms, i.e., algorithms that run in time $f(k)\cdot n^{\mathcal{O}(1)}$, for some computable function $f$. Such algorithms provide a fine-grained understanding on the time complexity of a problem and describe regions of the input space where the problem can be solved in polynomial time. Note that it is crucial here to obtain good bounds on the function $f$ since the (potentially) super-polynomial part of the running time is confined in the $f(k)$ term. In this direction it was proved that under the Exponential Time Hypothesis of Impagliazzo, Paturi and Zane, FVS does not admit an algorithm with running time $2^{o(n)}n^{\mathcal{O}(1)}$ (see [CFK+15]). Nevertheless certain classes of graphs (typically planar graphs) have been shown to admit algorithms with running times of the form $2^{\mathcal{O}(k^{\varepsilon})}n^{\mathcal{O}(1)}$ (for some $\varepsilon<1$), i.e., where the contribution of the parameter $k$ is subexponential. Such algorithms are called subexponential parameterized algorithms and they are the topic of this paper. Given the numerous existing results on this theme, there are two main directions of research: to improve the running times in the classes where an algorithm is already known, or to extend the tractability horizon of FVS by providing more general settings where subexponential FPT algorithms exist.
We are here interested by the second direction, that can be summarized by the following question.
@@
Historically, a primary source of graph classes studied to make progress on the above question was geometric intersection graphs. In an intersection graph, each vertex corresponds to a subset of some ambient space, and two vertices are adjacent if and only if the subsets intersect. Taking the Euclidean plane as the ambient space, many graph classes can been defined by setting restrictions on the subsets used to represent the vertices. One can for instance consider intersection graphs of disks in the plane, or segments, or Jordan arcs [^1] In the following, those are called strings.. With such subsets, one defines the class of disk graphs, segment graphs, and string graphs respectively. It is also often the case that there are conditions dealing with all the $n$ subsets representing the vertices of a given graph. For example, if we consider disks (resp. segments) one can ask those to have the same diameter (resp. to use at most $d$ different slopes), and this defines the class of unit disk graphs (resp. $d$-DIR graphs). When considering strings, one possible property is that any string has at most $d$ points shared with the other considered strings. This defines string graphs with edge degree at most $d$. A weaker condition is to ask every pair of the considered strings to intersect on at most $s$ points, this defines $s$-string graphs. This class generalizes several natural classes such as planar graphs, map graphs, unit-disk graphs, segment graphs, string graphs of bounded edge-degree, and intersection graphs of $\alpha$-convex bodies that exclude a fixed subgraph (see [Mat14b, KM94, BT22]).
@@
For all these graph classes, FVS is NP-complete, and actually under ETH none of them admits a $2^{o(\sqrt{n})}$-time algorithm. Indeed, this lower bound was given in [dBBKB+20] for induced grid graphs, which form a subclass of unit disk graphs and 2-DIR graphs. On the other hand for each of the aforementioned classes there is an algorithm solving FVS in subexponential time. More precisely, this algorithm applies to any string graph and runs in time $2^{\operatorname{\tilde{\mathcal{O}}}(n^{2/3})}$ [BR19]. [^2] The notation $\operatorname{\tilde{\mathcal{O}}}$ ignores polylogarithmic factors, i.e. we write $g(x)=\operatorname{\tilde{\mathcal{O}}}(f(x))$ if for some $c$ we have $g(x)=\mathcal{O}(f(x)\cdot\log^{c}x)$.
@@
Regarding subexponential parameterized algorithms, the case of unit disk graphs was settled with an algorithm whose running time matches the ETH lower bound [AO21]. This result uses the fact that these graphs admit vertex partitions into cliques such that each of these cliques is adjacent to only a constant number of the other cliques. Such a property does not hold for the other graph classes mentioned above. However, other techniques have been developed to deal with the other aforementioned classes such as the classes of bounded edge degree string graphs [BT22], contact-segment and square graphs [BBGR24a], disk graphs [LPS+22, ACO23], or the pseudo-disk graphs [BBGR25]. Note that when dealing with classes of intersection graphs, the representation of the input (if known) could be used by the algorithm. Some of these algorithms are robust, meaning that the input graph $G$ is provided using one of the classical graph data structures, where there is no indication of the intersection model of $G$. Because the recognition problem is difficult for most of the classes discussed above, robustness is a substantial advantage.
@@ 1.2. Our contribution
Let us now provide some intuition behind the conditions we require for a nice graph class. We discuss here the similarities between these conditions and classical studied properties, while the reasons why these conditions help to get a subexponential parameterized algorithm for FVS are examined in Section 2. A starting point is to review known results about string graphs, which constitute a good candidate to answer the previous question. In particular, the following results are known for string graphs. For a graph $H$, let us say that a graph is $H$-free if it does not contain $H$ as subgraph.
Theorem 1.2 ([Lee17],[DN19]).
@@
Theorem 1.3 ([Lee17]).
These results [^3] Note that an error was found in the proof of the above results in [Lee17], but a claim by the author was made that the proof can be corrected in the case of string graphs (see [BR24]). Moreover the earlier bound in [Mat14a] yields similar results, up to logarithmic factors. are interesting in our case, as a simple folklore branching allows us to reduce the problem to the case where the instance $(G,k)$ of FVS is $K_{r,r}$-free for $r=\lceil k^{\varepsilon}\rceil$. Thus, among the conditions required for a graph class to be nice, two of them correspond to a relaxed version of the above theorems.
Our last main condition is related to neighborhood complexity. A graph class $\mathcal{G}$ has linear neighborhood complexity (with ratio $c$) if for any graph $G\in\mathcal{G}$ and any $X\subseteq V(G)$, |$|\{N(v)\cap X,v\in V(G)\}|\leq c|X|$. It is known that bounded-expansion graph classes have linear neighborhood complexity [RVS19] as well as bounded twin-width graphs [BFLP24]. In previous work on parameterized subexponential algorithms [LPS+23, ACO23, BBGR25, BBGR24a], it appeared useful that the considered graphs have the property that, if $G$ is $K_{r,r}$-free (or even $K_{r}$-free), then for any $X\subseteq V(G)$, |$|\{N(v)\cap X,v\in V(G)\}|\leq r^{\mathcal{O}(1)}|X|$. Notice that this is slightly stronger than requiring that a class that is $K_{r,r}$-free (or $K_{r}$-free) for a fixed $r$ has linear neighborhood complexity, as it is important for our purpose that the dependency in $r$ is polynomial. We point out that $K_{r,r}$-free string graphs have bounded-expansion, hence linear neighborhood complexity, however this does not imply that the dependency in $r$ is polynomial. Thus, our last main condition (called bounded tree neighborhood complexity) can be seen as a slightly stronger version of this “polynomially dependent” neighborhood complexity. Let us now proceed to the formal definitions.
@@
We say that a graph class $\mathcal{G}$ has bounded tree neighborhood complexity (with parameters $\alpha,f_{1},f_{2}$) if there exist an integer $\alpha$ and two polynomial functions $f_{1},f_{2}$ such that the following conditions hold. For every $r$, every $K_{r,r}$-free graph $G\in\mathcal{G}$, every set $A\subseteq V(G)$ and every family $\mathcal{T}$ of disjoint non-adjacent [^4] Two vertex subsets are non-adjacent in a graph if there is no edge from one to the other, see Subsection 1.3. vertex subsets of $G-A$, each inducing a tree:
@@
    |$|\{N_{A}(T),~{}T\in\mathcal{T}\}|\leq f_{1}(r)|A|^{\alpha}$, where $N_{A}(T)$ denotes the neighbors of the vertices of $T$ in $A$, and
@@
    |$|\{N_{A}(T),~{}T\in\mathcal{T}\}|\leq f_{2}(r,p,m)|A|$, where $p$ and $m$ denote the maximum over all $T\in\mathcal{T}$ of $|N_{A}(T)|$ and $|T|$ respectively.
@@
We say that an hereditary graph class $\mathcal{G}$ is nice (for parameters $\alpha,f_{1},f_{2},\delta,f,d$) if all the following conditions hold:
  1. (1)
@@
  3. (3)
    There exist $\delta<1$ and a constant $f_{r}$ that depends polynomially in $r$ such that for any $K_{r,r}$-free graph $G\in\mathcal{G}$, $\operatorname{\textsf{tw}}(G)=\mathcal{O}(f_{r}\cdot n^{\delta})$.
  4. (4)
    There is a constant $d_{r}$ that depends polynomially in $r$ such that for any $K_{r,r}$-free graph $G\in\mathcal{G}$, $|E(G)|\leq d_{r}\cdot|V(G)|$. Without loss of generality we will assume $d_{r}\geq r$.
@@
Actually we provide a single generic algorithm for all nice classes and the parameters of the class (in the definition of nice) appear in the complexity analysis and are used to define $\eta$. The techniques used to prove the above result are discussed in Section 2. For the time being, let us focus on consequences. As hinted above, being nice is a natural property shared by several well-studied classes of graphs. In particular we show that it is the case for $s$-string graphs and pseudo-disk graphs, hence we have the following applications.
@@
There exists $\eta<1$, such that for all $s$ there is a robust parameterized subexponential algorithm solving FVS in time $2^{\operatorname{\tilde{\mathcal{O}}}\left(s^{\mathcal{O}(1)}k^{\eta}\right)}n^{\mathcal{O}(1)}$ for $n$-vertex $s$-string graphs.
@@
There exists $\eta<1$, such that there is a robust parameterized subexponential algorithm solving FVS in time $2^{k^{\eta}}n^{\mathcal{O}(1)}$ for $n$-vertex pseudo-disk graphs.
Observe that the two corollaries above encompass a wide range of classes of geometric intersection graphs for which subexponential parameterized algorithms have been given in previous work such as planar graphs, map graphs, unit-disk graphs, disk graphs, or more generally pseudo-disk graphs, and string graphs of bounded edge-degree. In this sense our main result unifies the previous algorithms.
Also, it captures new natural classes such as segment graphs, or more generally $s$-string graphs, where previous tools were unsuitable (as discussed in Section 2). We point out that before this work, the existence of subexponential parameterized algorithm for FVS was open even for the very restricted class of $2$-DIR graphs.
Generality has a cost and the running time bound we obtain for pseudo-disk graphs is worst than the one obtained in [BBGR25], which heavily relied on the input pseudo-disk representation. On the other hand our algorithm has the extra property of being robust (i.e., it does not require a geometric representation) which is a relevant advantage for those classes of intersection graphs where computing a representation is difficult.
1.3. Basic notations and organisation of the paper
In this paper logarithms are binary and all graphs are non-oriented and simple. Unless otherwise specified we use standard graph theory terminology, as in [Die05] for instance. For a graph $G$, and $v\in V(G)$, we denote $N_{G}(v)$ the neighbors of $v$. We omit the subscript when it is clear from the context. For $A\subseteq V(G)$, we use the notation $N(A)=\left(\cup_{v\in A}N(v)\right)\setminus A$ and denote $G[A]$ the subgraph induced by $G$ on $A$. For $v\in V(G)$ and $B\subseteq V(G)$, we denote $N_{B}(v)=N(v)\cap B$ and for $A\subseteq V(G)$ we denote $N_{B}(A)=N(A)\cap B$, with the additional notation $d_{B}(A)=|N_{B}(A)|$. For a graph $H$ we say that $G$ is $H$-free if $H$ is not a subgraph of $G$. Two disjoint vertex subsets or subgraphs $Z,Z^{\prime}$ of a graph $G$ are said to be non-adjacent (in $G$) if there is no edge in $G$ with one endpoint in $Z$ and the other in $Z^{\prime}$.
Organisation.
In Section 2 we explain why the approaches developed for other intersection graph classes in the papers [ACO23, BT22, BBGR25, LPS+22] do not apply here and present the main ideas behind our algorithm. Section 3 and Section 4 are devoted to the description and analysis of the algorithm. In Section 5 we provide applications of our main theorem by showing in particular that $s$-string graphs are nice. Finally, in Section 6 we discuss open problems and possible extensions of the approach developed here.

### v2

@@ Abstract.
The paper deals with the Feedback Vertex Set problem parameterized by the solution size. Given a graph $G$ and a parameter $k$, one has to decide if there is a set $S$ of at most $k$ vertices such that GSG-S is acyclic. Assuming the Exponential Time Hypothesis, it is known that FVS cannot be solved in time $2^{o(k)}n^{\mathcal{O}(1)}$ in general graphs. To overcome this, many recent results considered FVS restricted to particular intersection graph classes and provided such $2^{o(k)}n^{\mathcal{O}(1)}$ algorithms.
In this paper we provide generic conditions on a graph class for the existence of an algorithm solving FVS in subexponential FPT time, i.e. time $2^{k^{\varepsilon}}\mathop{\rm poly}(n)$, for some $\varepsilon<1$, where $n$ denotes the number of vertices of the instance and $k$ the parameter. On the one hand this result unifies algorithms that have been proposed over the years for several graph classes such as planar graphs, map graphs, unit-disk graphs, pseudo-disk graphs, and string graphs of bounded degree. On the other hand, it extends the tractability horizon of FVS to new classes that are not amenable to previously used techniques, in particular intersection graphs of “thin” objects like segment graphs or more generally $s$-string graphs.
@@ 1.1. Context
Given an $n$-vertex graph $G$ and a parameter $k\in\mathbb{N}$, the Feedback Vertex Set problem (FVS for short) asks whether there exists a set $S$ of at most $k$ vertices such that GSG-S has no cycle. This is a fundamental decision problem in graph theory and one of Karp’s 21 NP-complete problems. Because of its hardness in a classical setting, the problem has been widely studied within the realm of parameterized complexity. This line of research aims to investigate the existence of FPT algorithms, i.e., algorithms that run in time $f(k)\cdot n^{\mathcal{O}(1)}$, for some computable function $f$. Such algorithms provide a fine-grained understanding on the time complexity of a problem and describe regions of the input space where the problem can be solved in polynomial time. Note that it is crucial here to obtain good bounds on the function $f$ since the (potentially) super-polynomial part of the running time is confined in the $f(k)$ term. In this direction it was proved that under the Exponential Time Hypothesis of Impagliazzo, Paturi and Zane, FVS does not admit an algorithm with running time $2^{o(k)}n^{\mathcal{O}(1)}$ (see [Cygan2015Book]). Nevertheless certain classes of graphs (typically planar graphs) have been shown to admit algorithms with running times of the form $2^{\mathcal{O}(k^{\varepsilon})}n^{\mathcal{O}(1)}$ (for some $\varepsilon<1$), i.e., where the contribution of the parameter $k$ is subexponential. Such algorithms are called subexponential parameterized algorithms and they are the topic of this paper. Among the numerous existing results on this theme, there are two main directions of research: to improve the running times in the classes where an algorithm is already known, or to extend the tractability horizon of FVS by providing more general settings where subexponential FPT algorithms exist.
We are here interested by the second direction, which can be summarized by the following question.
@@
Historically, a primary source of graph classes studied to make progress on the above question was geometric intersection graphs. In an intersection graph, each vertex corresponds to a subset of some ambient space, and two vertices are adjacent if and only if the subsets intersect. Taking the Euclidean plane as the ambient space, many graph classes can be defined by setting restrictions on the subsets used to represent the vertices. One can for instance consider intersection graphs of disks in the plane, or segments, or Jordan arcs. [^1] In the following, those are called strings. With such subsets, one defines the class of disk graphs, segment graphs, and string graphs respectively. It is also often the case that there are conditions dealing with all the $n$ subsets representing the vertices of a given graph. For example, if we consider disks (resp. segments) one can ask those to have the same diameter (resp. to use at most $d$ different slopes), and this defines the class of unit disk graphs (resp. $d$-DIR graphs). When considering strings, one possible property is that any string has at most $d$ points shared with the other considered strings. This defines string graphs with edge-degree at most $d$. It was recently proved that for a string graph, having bounded edge-degree is in fact equivalent to having bounded degree [karol2025]. A weaker condition is to ask every pair of the considered strings to intersect on at most $s$ points, this defines $s$-string graphs. This class generalizes several natural classes such as planar graphs, map graphs, unit-disk graphs, segment graphs, string graphs of bounded degree, and intersection graphs of $\alpha$-convex bodies that exclude a fixed subgraph (see [matouvsek2014string, kratochvil1994intersection, baste2022contraction]).
@@
For all these graph classes, FVS is NP-complete, and actually under ETH none of them admits a $2^{o(\sqrt{n})}$-time algorithm. Indeed, this lower bound was given in [deberg_unit] for induced grid graphs, which form a subclass of unit disk graphs and 2-DIR graphs. On the other hand, for each of the aforementioned classes there is an algorithm solving FVS in subexponential time. More precisely, this algorithm applies to any string graph and runs in time $2^{\operatorname{\tilde{\mathcal{O}}}(n^{2/3})}$ [bonnet2019optimality]. [^2] The notation $\operatorname{\tilde{\mathcal{O}}}$ ignores polylogarithmic factors, i.e. we write $g(x)=\operatorname{\tilde{\mathcal{O}}}(f(x))$ if for some $c$ we have $g(x)=\mathcal{O}(f(x)\cdot\log^{c}x)$.
@@
Regarding subexponential parameterized algorithms, the case of unit disk graphs was settled with an algorithm whose running time matches the ETH lower bound [an_unit_21]. This result uses the fact that these graphs admit vertex partitions into cliques such that each of these cliques is adjacent to only a constant number of the other cliques. Such a property does not hold for the other graph classes mentioned above. However, other techniques have been developed to deal with the other aforementioned classes such as the classes of bounded edge-degree string graphs [baste2022contraction], contact-segment and square graphs [berthe24ASQGM], disk graphs [lokSODA22, Faster2023Shinwoo], or the pseudo-disk graphs [FVS-WG]. Note that when dealing with classes of intersection graphs, the representation of the input (if known) could be used by the algorithm. Some of these algorithms are robust, meaning that the input graph $G$ is provided using one of the classical graph data structures, where there is no indication of the intersection model of $G$. Because the recognition problem is difficult for most of the classes discussed above, robustness is a substantial advantage.
@@ 1.2. Our contribution
Let us now provide some intuition behind the conditions we require for a nice graph class. We discuss here the similarities between these conditions and classical studied properties, while the reasons why these conditions help to get a subexponential parameterized algorithm for FVS are examined in Section 2. A starting point is to review known results about the class of string graphs, which constitutes a good candidate to answer the previous question. In particular, the following results are known for string graphs. For a graph $H$, let us say that a graph is $H$-free if it does not contain $H$ as a subgraph.
Theorem 1.2 ([lee2016separators]).
Theorem 1.3 ([lee2016separators, DVORAK2019137]).
@@
These results [^3] Note that an error was found in the proof of the above results in [lee2016separators], but a claim by the author was made that the proof can be corrected in the case of string graphs (see [Bonnet2024]). Moreover the earlier bound in [Matousek14] yields similar results, up to logarithmic factors. are interesting in our case, as a simple folklore branching allows us to reduce the problem to the case where the instance $(G,k)$ of FVS is $K_{r,r}$-free for $r=\lceil k^{\varepsilon}\rceil$. Thus, among the conditions required for a graph class to be nice, two of them correspond to a relaxed version of the above theorems.
Our last main condition is related to neighborhood complexity. A graph class $\mathcal{G}$ has linear neighborhood complexity (with ratio $c$) if for any graph $G\in\mathcal{G}$ and any $X\subseteq V(G)$, we have that $|\{N(v)\cap X,v\in V(G)\}|\leq c|X|$. It is known that graph classes of bounded expansion have linear neighborhood complexity [reidl2019characterising] as well as graph classes of bounded twin-width [bonnet2024neighbourhood]. In previous works on parameterized subexponential algorithms [Lokshtanov23Approx, Faster2023Shinwoo, FVS-WG, berthe24ASQGM], it appeared useful that the considered graphs have the property that, if $G$ is $K_{r,r}$-free (or even $K_{r}$-free), then for any $X\subseteq V(G)$, we have that $|\{N(v)\cap X,v\in V(G)\}|\leq r^{\mathcal{O}(1)}|X|$. Notice that this is slightly stronger than requiring that a class that is $K_{r,r}$-free (or $K_{r}$-free) for a fixed $r$ has linear neighborhood complexity, as it is important for our purpose that the dependency in $r$ is polynomial. We point out that $K_{r,r}$-free string graphs have bounded-expansion (by Theorem 1.3 and [Esperet2018]), hence linear neighborhood complexity, however this does not imply that the dependency in $r$ is polynomial. Thus, our last main condition (called bounded tree neighborhood complexity) can be seen as a slightly stronger version of this “polynomially dependent” neighborhood complexity. Let us now proceed to the formal definitions.
@@
We say that a graph class $\mathcal{G}$ has bounded tree neighborhood complexity (with parameters $\alpha,f_{1},f_{2}$) if there exist an integer $\alpha$ and two polynomial functions $f_{1},f_{2}$ such that the following conditions hold. For every $r$, every $K_{r,r}$-free graph $G\in\mathcal{G}$, every set $A\subseteq V(G)$ and every family $\mathcal{T}$ of disjoint non-adjacent [^4] Two vertex subsets are non-adjacent in a graph if there is no edge from one to the other, see Subsection 1.3. vertex subsets of GAG-A, each inducing a tree:
@@
    $|\{N_{A}(T),~T\in\mathcal{T}\}|\leq f_{1}(r)|A|^{\alpha}$, where $N_{A}(T)$ denotes the neighbors of the vertices of $T$ in $A$, and
@@
    $|\{N_{A}(T),~T\in\mathcal{T}\}|\leq f_{2}(r,p,m)|A|$, where $p$ and $m$ denote the maximum over all $T\in\mathcal{T}$ of $|N_{A}(T)|$ and $|T|$ respectively.
@@
We say that an hereditary graph class $\mathcal{G}$ is nice (for parameters $\alpha,f_{1},f_{2},\delta,\tau,d$) if all the following conditions hold:
  1. (C1)
@@
  3. (C3)
    There exist a constant $\delta<1$ and a polynomial function $\tau(r)$ such that for any $K_{r,r}$-free graph $G\in\mathcal{G}$, $\operatorname{\textsf{tw}}(G)\leq\tau(r)\cdot n^{\delta}$.
  4. (C4)
    The density $|E(G)|/|V(G)|$, of any $K_{r,r}$-free graph $G\in\mathcal{G}$, is upper bounded by a polynomial function $d(r)$. Without loss of generality we will assume $d(r)\geq r$.
@@
For every nice hereditary graph class $\mathcal{G}$ there is a constant $\eta<1$ such that FVS can be solved in $\mathcal{G}$ in time $2^{\mathcal{O}(k^{\eta})}\cdot n^{\mathcal{O}(1)}$.
Actually, we describe a single algorithm parameterized by the parameters of the nice classes it applies to. These parameters are used to make choices during the execution of the algorithm and they are used to define $\eta$. The techniques used to prove the above result are discussed in Section 2. For the time being, let us focus on consequences. As hinted above, being nice is a natural property shared by several well-studied classes of graphs. In particular we show that it is the case for $s$-string graphs and pseudo-disk graphs, hence we have the following applications.
@@
There exists $\eta<1$, such that for all $s$ there is a robust algorithm solving FVS in time $2^{\operatorname{\tilde{\mathcal{O}}}\left(s^{\mathcal{O}(1)}k^{\eta}\right)}n^{\mathcal{O}(1)}$ for $n$-vertex $s$-string graphs.
@@
There exists $\eta<1$, such that there is a robust algorithm solving FVS in time $2^{\mathcal{O}(k^{\eta})}n^{\mathcal{O}(1)}$ for $n$-vertex pseudo-disk graphs.
Observe that the two corollaries above encompass a wide range of classes of geometric intersection graphs for which subexponential parameterized algorithms have been given in previous works such as planar graphs, map graphs, unit-disk graphs, disk graphs, or more generally pseudo-disk graphs, and string graphs of bounded degree. In this sense our main result unifies the previous algorithms.
Also, it captures new natural classes such as segment graphs, or more generally $s$-string graphs, where previous tools were not sufficient (as discussed in Section 2). We point out that before this work, the existence of subexponential parameterized algorithm for FVS was open even for the very restricted class of $2$-DIR graphs.
1.3. Basic notation
In this paper logarithms are binary and all graphs are non-oriented and simple. Unless otherwise specified we use standard graph theory terminology, as in [diestel2005graph] for instance. For a graph $G$, and $v\in V(G)$, we denote $N_{G}(v)$ the neighbors of $v$. We omit the subscript when it is clear from the context. For $A\subseteq V(G)$, we use the notation $N(A)=\left(\cup_{v\in A}N(v)\right)\setminus A$ and denote $G[A]$ the subgraph induced by $G$ on $A$. For $v\in V(G)$ and $B\subseteq V(G)$, we denote $N_{B}(v)=N(v)\cap B$ and for $A\subseteq V(G)$ we denote $N_{B}(A)=N(A)\cap B$, with the additional notation $d_{B}(A)=|N_{B}(A)|$. For a graph $H$ we say that $G$ is $H$-free if $H$ is not a subgraph of $G$. Two disjoint vertex subsets or subgraphs $Z,Z^{\prime}$ of a graph $G$ are said to be non-adjacent (in $G$) if there is no edge in $G$ with one endpoint in $Z$ and the other in $Z^{\prime}$.
1.4. Organisation of the paper
In Section 2 we explain why the approaches developed for other intersection graph classes in the papers [Faster2023Shinwoo, baste2022contraction, FVS-WG, lokSODA22] do not apply here and present the main ideas behind our algorithm. Section 3 and Section 4 are devoted to the description and analysis of the algorithm. In Section 5 we provide applications of our main theorem by showing in particular that $s$-string graphs are nice. Finally, in Section 6 we discuss open problems and possible extensions of the approach developed here.